OpenStudy (skittles_for_life6422):

!!!!!!!!!!!!!!!!!!!!!!!!!WILL GIVE MEDAL AND FAN!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! What are the domain and ranges of the function

OpenStudy (skittles_for_life6422):

\[3\sqrt{6x+42}\]

OpenStudy (skittles_for_life6422):

@StudyGurl14

OpenStudy (skittles_for_life6422):

hi @StudyGurl14

OpenStudy (studygurl14):

Hi @skittles_for_life6422 :)

OpenStudy (skittles_for_life6422):

can you help with a couple problems

OpenStudy (skittles_for_life6422):

i think the answer for this is \[x \ge-7\]

OpenStudy (studygurl14):

I'm not sure actually...for some reason I can't remember how to find domain for a question with square root...

OpenStudy (skittles_for_life6422):

thats ok

OpenStudy (skittles_for_life6422):

@BIGDADDY69420 CAN YOU HELP WITH THIS QUESTION

OpenStudy (skittles_for_life6422):

@jdoe0001

OpenStudy (skittles_for_life6422):

@SolomonZelman

OpenStudy (skittles_for_life6422):

@TheSmartOne can you help with my question

OpenStudy (skittles_for_life6422):

@bohotness can you help with my question

OpenStudy (anonymous):

whats the question?

OpenStudy (skittles_for_life6422):

first thing one this chat

OpenStudy (skittles_for_life6422):

wait that one is the wrong one this the right one \[2\sqrt{3x+4-5}\]

OpenStudy (skittles_for_life6422):

@TheSmartOne

OpenStudy (anonymous):

what are the options?